OpenStudy (anonymous):

Find the slope of a line 4x-7y=14

OpenStudy (aroub):

-7y=-4x+14 y=4/7x-2 slope=4/7

OpenStudy (amistre64):

4x - 7y = 14 /-7 /-7 /-7 -------------- (-4/7)x + y = -2 +(4/7)x +(4/7)x ------------------- y = (4/7)x -2